OpenStudy (anonymous):

A field is 500 m long and 30 m broad. A ditch 30 m long, 20 m broad and 12 m deep is dug in the field and the earth taken out is evenly spread over the field. Find the level of the field raised.

OpenStudy (anonymous):

First, figure out the volume of the dirt dug up: LxBxD = 30x20x12 = 7200 cubic meters Then, figure out the surface area of the "undug" field. This would be the whole field, minus the surface size of the hole: (500x30) - (30x20) = 15000 -600 = 14400 sq meters So now we know that 7200 cubic meters of dirt will be spread over 14400 square meters of land. (Are you making a mental picture of this - I'm picturing cubes of dirt - I don't have enough to put one on each square meter of land, so I know my answer will be a fraction...) Divide the volume of dirt by the square footage it has to cover. 7200/14400 = 1/2 meter of dirt, or .5 meter (or 500 cm...)

OpenStudy (anonymous):

yeahhh ..... u r totally correct but let me tell u that 1/2 meter is 50 cm.....not 500 cm...:DDD.....lol